OpenStudy (anonymous):

Hiram Callaway deposits $2,350 in an account that earns interest at 3.2% APR compounded monthly. What is his ending balance after five years?

jimthompson5910 (jim_thompson5910):

Formula \[\Large F = P*\left(1+\frac{r}{n}\right)^{n*t}\] F = final value P = present value r = interest rate as a percentage n = number of times money is compounded per year t = number of years

OpenStudy (anonymous):

Your problem would be \[F = 2350(1+\frac{ 0.032 }{ 12})^{12*5}\] I believe

OpenStudy (wolf1728):

Total = 2,350 * (1+ [.032/12]) ^ 60 Total = 2,350 * (1.0026666667)^60 Total = 2,350 * 1.1732609928 Total = 2,757.16 and a calculator to check that answer: http://www.1728.org/compint.htm